OpenStudy (anonymous):

f(x) = 1/3x − 2 and g(x) = −2x − 1, find (f • g)(9)

OpenStudy (anonymous):

f(g(9))=f(-2(9)-1)=f(-18-1)=f(-19).... put x=-19 in f(x)

OpenStudy (stamp):

Find g(9) and plug it in for f(x)\[f(g(9))=\frac{1}{3}g(9)-2\]

OpenStudy (anonymous):

-19?